OpenStudy (raffle_snaffle):

A company has annual fixed costs of $2983007 and variable costs of $0.73 per unit produced. If the company charges $1.88 per unit for their product, what is the breakeven annual production quantity? Assume that the demand is unlimited.

OpenStudy (raffle_snaffle):

@DanJS

OpenStudy (raffle_snaffle):

@DanJS

OpenStudy (raffle_snaffle):

Total cost = total fixed cost + total variable cost

OpenStudy (danjs):

looks like you get a cost function, and a revenue function, let x be the units produced C(x) = 0.73*x + 2983007 R(x) = 1.88*x what is the break even thing, just have the cost and the revenue the same value?

OpenStudy (danjs):

amount of units to produce to break even i guess

OpenStudy (raffle_snaffle):

Can you please show me or we can discuss how we analyze this type of problem?

OpenStudy (raffle_snaffle):

yes I believe you are correct. Are we using the idea of the equation i posted?

OpenStudy (danjs):

looks like you understood the idea for the cost function and revenue total $ = initial amt or fixed value + number of something*value

OpenStudy (danjs):

the cost is the fixed value plus the cost of each unit made

OpenStudy (danjs):

revenue is just the sale price times how many sold here solve for units x, when C(x) = R(x)

OpenStudy (raffle_snaffle):

So the variable cost is the change in the unit cost?, hence varible?

OpenStudy (raffle_snaffle):

total cost = total revenue?

OpenStudy (danjs):

just a fancy word, the cost there changes depending on how many units are made, so it is variable i guess

OpenStudy (raffle_snaffle):

In my textbook they use "total variable cost"

OpenStudy (raffle_snaffle):

so total cost is given by C(x)?

OpenStudy (danjs):

what is your book meaning for 'breakeven annual production quantity', i am guessing just figure the amount of units x, that will have cost = revenue for the whole year

OpenStudy (danjs):

0.73*x + 2983007 = 1.88x

OpenStudy (raffle_snaffle):

text says break even point: The level of activity at which total cost for the product, good, or service are equal to the revenue generated. This is the level at which one "just breaks even" The math part I get... I want to understand how the solution part of it.

OpenStudy (raffle_snaffle):

Thanks for your help.

OpenStudy (raffle_snaffle):

x=2.59x10^6

OpenStudy (danjs):

right, break even, cost = revenue, no profits

OpenStudy (danjs):

welcome